# Exploring the effect of different cosmologies on the Epoch of Reionization 21-cm signal with Polar

### v1

@@ Appendix A UV Luminosity Functions across redshifts
Figure 9: As Figure 2, but for the constrained case UVLFs at $z=12$, 10, 9, 8, 7 and 5.
Appendix B Maps of $\delta T_{\rm b}$
In Figure 10, we show the maps of $\delta T_{\rm b}$ of the B-series middle slices (the simulations with “paired” initial conditions in the F&P pair of simulations) for the four cosmological models in the constrained case at $z=12,10,8,~{}\rm and~{}6$ (from top to bottom). Here the darkest regions represent the ionized regions with $\delta T_{\rm b}=0$. Note that $\delta T_{\rm b}$ cannot assume negative values, due to the assumption of $T_{\rm S}\gg T_{\rm CMB}$. We note that the ionized regions of the A-series middle slices shown in Figure 3, correspond to the most neutral regions of Figure 10. This is because of the nature of the Fixed & Paired approach, as regions of matter clustering in the A-series should correspond to voids in the B-series by construction.
Figure 10: Same as Figure 3 but for the B-series of simulations.
Appendix C Fiducial model with $\text{f}_{\text{esc}}=25\%$
@@
Figure 11: As Figure 4, but for the fiducial model, for an escape fraction of $f_{\rm esc}=12.5\%$ (cyan) and 25% (brown).
Figure 12: As Figure 8, but for the fiducial model, for an escape fraction of $f_{\rm esc}=12.5\%$ (cyan) and 25% (brown).

### v2

@@ Appendix A UV Luminosity Functions across redshifts
Figure 9: A and B-series averaged UVLFs for the constrained case at $z=12$, 10, 9, 8, 7 and 5 for the fiducial (cyan), $h$ high (red), $\sigma_{8}$ low (purple) and $\sigma_{8}$ high (green). We also show the dust attenuated (solid) and unattenuated (dashed) UVLFs, and JWST and HST observations (Finkelstein et al., 2015; Bouwens et al., 2015, 2021; Harikane et al., 2022; Bouwens et al., 2023a; Bouwens et al., 2023b; Harikane et al., 2023; Leung et al., 2023; McLeod et al., 2024; Adams et al., 2024, grey circles).
Appendix B Fiducial model with $\text{f}_{\text{esc}}=25\%$
@@
Figure 10: Redshift evolution of the average of the A and B-series volume-averaged neutral hydrogen fraction $\langle x_{\rm HI}\rangle$ for the fiducial model, for an escape fraction of $f_{\rm esc}=12.5\%$ (cyan) and 25% (brown). Dotted lines refer to a fit to the curves, which is used for a better estimate of the redshift of reionization when the redshift resolution is too coarse. The vertical grey dashed lines indicate the redshifts observationally relevant for LOFAR ($z=10.11$, 9.16 and 8.3), and the black solid line is drawn at $\langle x_{\rm HI}\rangle=0.5$ to guide the eye. Grey circles are a collection of observational constraints (Fan et al., 2006; Totani et al., 2006; Ota et al., 2008; Ouchi et al., 2010; Bolton et al., 2011; Dijkstra et al., 2011; McGreer et al., 2011; Mortlock et al., 2011; Ono et al., 2012; Chornock et al., 2013; Jensen et al., 2013; Robertson et al., 2013; Schroeder et al., 2013; Pentericci et al., 2014; Schenker et al., 2014; McGreer et al., 2015; Sobacchi & Mesinger, 2015; Choudhury et al., 2015; Mesinger et al., 2015; Greig et al., 2017; Davies et al., 2018; Mason et al., 2018; Hoag et al., 2019; Greig et al., 2019; Jones et al., 2024, collected in the CoReCon module, Garaldi 2023).
Figure 11: Redshift evolution of the A and B-series averaged normalized 21-cm signal power spectrum ($\Delta^{2}_{\rm 21cm}$) at $k=0.15~h\rm cMpc^{-1}$ for the fiducial model, for an escape fraction of $f_{\rm esc}=12.5\%$ (cyan) and 25% (brown). The vertical dashed lines indicate the redshifts relevant for LOFAR, i.e. $z=10.11$, 9.16, and 8.3.
Appendix C Maps of $\delta T_{\rm b}$
C.1 B-series maps in the constrained case
In Figure 12, we show the maps of $\delta T_{\rm b}$ of the B-series middle slices (the simulations with “paired” initial conditions in the F&P pair of simulations) for the four cosmological models in the constrained case at $z=12,10,8,~\rm and~6$ (from top to bottom). Here the darkest regions represent the ionized regions with $\delta T_{\rm b}=0$. Note that $\delta T_{\rm b}$ cannot assume negative values, due to the assumption of $T_{\rm S}\gg T_{\rm CMB}$. We note that the ionized regions of the A-series middle slices shown in Figure 3 correspond to the most neutral regions of Figure 12. This is because of the nature of the Fixed & Paired approach, as regions of matter clustering in the A-series should correspond to voids in the B-series by construction.
Figure 12: Maps of $\delta T_{\rm b}$ of the B-series middle slices of single cell thickness (i.e., $\approx 600~h^{-1}~\rm ckpc$) for the four cosmological models in the constrained case at $z=12,10,8,~\rm and~6$ (from top to bottom) corresponding to the A-series slices shown in Figure 3. Here the dark areas represent the ionized regions with $\delta T_{\rm b}=0$. Note that $\delta T_{\rm b}$ cannot have negative values due to the assumption of $T_{\rm S}\gg T_{\rm CMB}$.
In Figure 13, we show the absolute difference between the maps of $\delta T_{\rm b}$ of the A and B-series middle slices for the four cosmological models in the constrained case at $z=12,10,8,~\rm and~6$ (from top to bottom). This makes the difference between the fixed simulation and its pair evident, as well as showing that with the progress of reionization, small-scale stochasticity induced by ionized bubbles is a tracer for deviations from the perfect anti-symmetry that the system started with. That is, until $z=8$, the middle slices are effectively random noise maps, as they are primarily governed by the underlying matter distribution. However, by $z=6$, the impact of the growth of ionized bubbles becomes significant and shows up as features in the difference maps.
Figure 13: The absolute difference between the maps of $\delta T_{\rm b}$ of the A and B-series middle slices shown in Figure 3 and 12.
C.2 A and B-series maps in the unconstrained case
In Figure 14 and 15, we show the maps of $\delta T_{\rm b}$ of the A and B-series middle slices for the four cosmological models in the unconstrained case at $z=12,10,8,~\rm and~6$ (from top to bottom). Thus, here, we directly see the influence of the choice of cosmological parameters on the 21-cm signal.
Figure 14: Maps of $\delta T_{\rm b}$ of the A-series middle slices of single cell thickness (i.e., $\approx 600~h^{-1}~\rm ckpc$) for the four cosmological models in the unconstrained case at $z=12,10,8,~\rm and~6$ (from top to bottom). Note that here, due to the higher matter clustering driving up star formation, the $\sigma_{8}$ high model is already fully ionized by $z=6$, and the $h$ high model is approaching complete reionization as well. On the other hand, the $\sigma_{8}$ low case still has large neutral regions.
Figure 15: Maps of $\delta T_{\rm b}$ of the B-series middle slices of single cell thickness (i.e., $\approx 600~h^{-1}~\rm ckpc$) for the four cosmological models in the unconstrained case at $z=12,10,8,~\rm and~6$ (from top to bottom).